# Convergence of Stochastic Gradient Langevin Dynamics in the Lazy Training Regime

### v1

@@ 3 Convergence of Stochastic Gradient Langevin Dynamics in the Lazy Training Regime
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau=\infty\}}]\xrightarrow[]{t\rightarrow\infty}0.$
We will return to this problem in Theorem 3 to control the error stemming from $\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<\infty\}}]$.
@@
where we only train the weights $\omega$ and randomly choose $c$ at initialization. This neural network fulfills the assumptions of Theorem 1. This implies that, for appropriate loss functions that fulfill Assumption 1 (such as the mean squared loss function), equation 3 holds for shallow neural networks.
@@
The proof is presented in Appendix 6.1. ∎
@@
The proof of Corollary 2 can be found in Appendix 6.2. ∎
@@
However, using the Lemmas B1-B4 in Du et al. (2019a), we can define a different first-exit time that guarantees the positive definiteness of the NTK. Instead of using the radius $r:=\lambda/\text{Lip}(Dh)$, we will use a different radius, for which the above proof can be conducted identically. This circumvents the need for $Dh$ to be Lipschitz-smooth. Further discussion is provided in Appendix 6.2.
@@
$\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|^{2}_{2}]\leq\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]\exp(-2m\lambda^{2}t).$
@@
$\frac{m}{2}\|\alpha h(\omega_{t})-h^{\star}\|_{2}^{2}\leq\bar{R}(\alpha h(\omega_{t})).$
@@
$\frac{m}{2}\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}]$ $\leq\mathbb{E}[\bar{R}(\alpha h(\omega_{0}))]\exp(-2m\lambda^{2}t)$
$\leq\frac{\mathrm{Lip}(\nabla R)}{2}\mathbb{E}[\|\alpha h(\omega_{0})-h^{\star}\|_{\mathbb{F}}^{2}]\exp(-2m\lambda^{2}t),$
where we used the smoothness inequality $R(\alpha h(\omega_{0}))-R(\alpha h(\omega^{\star}))\leq\frac{1}{2}\mathrm{Lip}(\nabla R)\|\alpha h(\omega_{0})-h(\omega^{\star})\|_{\mathbb{F}}^{2}$ by Lemma 3.4 in Bubeck et al. (2015). This leads to the desired equation
$\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}^{2}]\leq\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}[\|\alpha h(\omega_{0})-h^{\star}\|_{2}^{2}]\exp(-2m\lambda^{2}t).$
We use the above results to show that by choosing $\alpha$ sufficiently large, we can ensure that $\tau=\infty$ with high probability since $\lambda Id\preceq Dh(\omega_{t})D^{T}h(\omega_{t})$ holds. In the following result and throughout the rest of the paper, we denote the conditional probability and expectation given the random initialization $\omega_{0}$ as $\mathbb{P}$ and $\mathbb{E}$.
@@
Proof of Theorem 2.
To prove the above statement, we will first use equation 1 and rewrite it to get an upper bound on the expected distance of $\omega_{t}$ from initialization. The resulting terms can then be further bounded by using the regularity of the loss function from Assumption 1 and the results from Theorem 1. We connect the expected distance to the probability of exiting a ball around the initialization using Markov’s inequality by $\mathbb{P}(\|\omega_{t}-\omega_{0}\|>\epsilon)\leq\frac{\mathbb{E}[\|\omega_{t}-\omega_{0}\|]}{\epsilon}$. Using equation 1, we can rewrite the distance of $\omega_{t}$ to the origin
$\leq\frac{1}{\alpha}\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{F}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt+\frac{\sqrt{\eta_{\alpha}}}{\alpha}\|\int_{0}^{t}\sigma_{\alpha}(\omega_{t})dB_{t}\|_{2}$
$\mathbb{E}[\|\omega_{t}-\omega_{0}\|]$ $\leq\frac{\mathbb{E}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{F}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{\alpha}+\frac{\sqrt{\eta_{\alpha}}\mathbb{E}[\|\int_{0}^{t}\sigma_{\alpha}(\omega_{t})dB_{t}\|_{2}]}{\alpha}$
$\leq\frac{\mathbb{E}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{F}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{\alpha}+\frac{\left[\mathbb{E}[\int_{0}^{t}\mathrm{Tr}(\Sigma_{\alpha}(\omega_{s})ds)]\right]^{\frac{1}{2}}}{\alpha}.$
From the proof of Theorem 1, we already know that $\int_{0}^{t}\mathrm{Tr}(\Sigma_{\alpha}(\omega_{s}))ds\leq 0$. Therefore, it follows that
$\mathbb{P}(\|\omega_{t}-\omega_{0}\|_{2}>r)$ $\leq\frac{\mathbb{E}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{F}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{r\alpha}$
$\overset{t\leq\tilde{T}}{\leq}\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\mathbb{E}[\int_{0}^{t}\|\alpha h(\omega_{t})-h^{\star}\|_{2}dt]$
$=\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\int_{0}^{t}\sqrt{\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}]^{2}}dt$
$\overset{(\spadesuit)}{\leq}\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\int_{0}^{t}\sqrt{\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}^{2}]}dt$
$\overset{(\clubsuit)}{\leq}\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\int_{0}^{t}\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]\exp(-2m\lambda^{2}t)}dt$
$=\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]}\int_{0}^{t}\exp(-m\lambda^{2}t)dt$
$\leq\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\frac{\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]}}{m^{\frac{3}{2}}\lambda^{2}},$
where $(\spadesuit)$ follows from Jensen’s inequality and $(\clubsuit)$ follows from Corollary 3. By choosing an initialization such that $h(\omega_{0})=0$, we then get
$\mathbb{P}(\|\omega_{t}-\omega_{0}\|_{2}>r)\leq\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\frac{\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}[\|h^{\star}\|^{2}_{2}]}}{m^{\frac{3}{2}}\lambda^{2}}.$
In the following corollary, we will show that by choosing $\alpha$ appropriately, $\mathbb{P}(\tau<\infty)$ is arbitrarily small.
Corollary 4.
Under the same assumptions as Theorem 2, it follows that
$\mathbb{P}(\tau<\infty)\leq\mathcal{O}(\alpha^{-1}).$ (10)
Proof of Corollary 4.
Since the right-hand side of equation 9 does not depend on $t$, we get
[…]
$\leq\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\frac{\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}[\|h^{\star}\|^{2}_{2}]}}{m^{\frac{3}{2}}\lambda^{2}}=\mathcal{O}(\alpha^{-1})$
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t}))]\leq\exp(-2\lambda^{2}mt)+\mathcal{O}(\alpha^{-\frac{1}{q}}),$
for any $q>1$.
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t}))]=\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau=\infty\}}]+\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<\infty\}}].$
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau=\infty\}}]\leq\exp(-2\lambda^{2}mt),$
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<\infty\}}]\leq\mathcal{O}(\alpha^{-\frac{1}{q}}).$
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<\infty\}}]$ $\leq\mathbb{E}[\bar{R}(\alpha h(\omega_{t}))^{p}]^{\frac{1}{p}}\mathbb{E}[\mathds{1}_{\{\tau<\infty\}}^{q}]^{\frac{1}{q}}$ (11)
$\overset{\spadesuit}{=}\mathbb{E}[\bar{R}(\alpha h(\omega_{t})^{p}]^{\frac{1}{p}}\mathbb{E}[\mathds{1}_{\{\tau<\infty\}}]^{\frac{1}{q}}$ (12)
$=\mathbb{E}[\bar{R}(\alpha h(\omega_{t})^{p}]^{\frac{1}{p}}\mathbb{P}(\tau<\infty)^{\frac{1}{q}}$ (13)
$\overset{\clubsuit}{\leq}\mathbb{E}[\bar{R}(\alpha h(\omega_{t})^{p}]^{\frac{1}{p}}\mathcal{O}(\alpha^{-\frac{1}{q}}),$ (14)
@@
$\mathbb{E}[\bar{R}(\alpha h(\omega_{t}))^{p}]\leq\bar{R}(\alpha h(\omega_{0}))^{p}e^{p}\mathbb{E}[\mathcal{E}_{t}]=\bar{R}(\alpha h(\omega_{0}))^{p}e^{p}.$
Plugging this result into equation 14 concludes the proofn. ∎
@@
$\mathbb{E}[\|\alpha\bar{h}(\bar{\omega}_{t})-h^{\star}\|_{2}^{2}]\leq\frac{\mathrm{Lip}(\nabla R)}{m}\|h^{\star}\|^{2}_{2}\exp(-2m\lambda^{2}t),$ (15)
@@
$\mathbb{P}(\|\bar{\omega}_{t}-\omega_{0}\|_{2}>r)\leq\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{F}\mathrm{Lip}(\nabla R)\frac{\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}[\|h^{\star}\|^{2}_{2}]}}{m^{\frac{3}{2}}\lambda^{2}}=\mathcal{O}\left(\alpha^{-1}\right).$
@@
$\mathbb{E}[\|\alpha\bar{h}(\bar{\omega}_{t})-\alpha h(\omega_{t})\|_{2}]\leq 2\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}}\|h^{\star}\|_{2}\exp(-m\lambda^{2}t).$
@@
$\mathbb{E}[\|\alpha\bar{h}(\bar{\omega}_{t})-\alpha h(\omega_{t})\|_{2}]$ $\leq\mathbb{E}[\|\alpha\bar{h}(\bar{\omega}_{t})-h^{\star}\|_{2}]+\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}]$
$\overset{(\clubsuit)}{\leq}\sqrt{\mathbb{E}[\|\alpha\bar{h}(\bar{\omega}_{t})-h^{\star}\|_{2}^{2}]}+\sqrt{\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}^{2}]}$
$\overset{(\spadesuit)}{\leq}2\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}}\|h^{\star}\|_{2}\exp(-m\lambda^{2}t),$

### v2

@@ 3 Convergence of Stochastic Gradient Langevin Dynamics in the Lazy Training Regime
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau=\infty\}}]\xrightarrow[]{t\rightarrow\infty}0.$
We will return to this problem in Theorem 3 to control the error stemming from $\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<\infty\}}]$.
@@
where we only train the weights $\omega$ and randomly choose $c$ at initialization. Here, $\sigma$ is the non-linear activation function $\sigma=\tanh$. This neural network fulfills the assumptions of Theorem 1. This implies that, for appropriate loss functions that fulfill Assumption 1 (such as the mean squared loss function), equation 3 holds for shallow neural networks.
@@
The proof is presented in Appendix B. ∎
@@
The proof of Corollary 2 can be found in Appendix C. ∎
@@
However, using the Lemmas B1-B4 in Du et al. (2019a), we can define a different first-exit time that guarantees the positive definiteness of the NTK. Instead of using the radius $r:=\lambda/\text{Lip}(Dh)$, we will use a different radius, for which the above proof can be conducted identically. This circumvents the need for $Dh$ to be Lipschitz-smooth. Further discussion is provided in Appendix C.
@@
$\mathbb{E}_{\omega}[\|\alpha h(\omega_{t})-h^{\star}\|^{2}_{\mathbb{F}}]\leq\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{\mathbb{F}}]\exp(-m\lambda^{2}t).$
@@
$\frac{m}{2}\|\alpha h(\omega_{t})-h^{\star}\|_{\mathbb{F}}^{2}\leq\bar{R}(\alpha h(\omega_{t})).$
@@
$\frac{m}{2}\mathbb{E}_{\omega}[\|\alpha h(\omega_{t})-h^{\star}\|_{\mathbb{F}}]$ $\leq\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{0}))]\exp(-m\lambda^{2}t)$
$\leq\frac{\mathrm{Lip}(\nabla R)}{2}\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|_{\mathbb{F}}^{2}]\exp(-m\lambda^{2}t),$
where we used the smoothness inequality $R(\alpha h(\omega_{0}))-R(h^{\star})\leq\frac{1}{2}\mathrm{Lip}(\nabla R)\|\alpha h(\omega_{0})-h^{\star}\|_{\mathbb{F}}^{2}$ by Lemma 3.4 in Bubeck et al. (2015). This leads to the desired equation
$\mathbb{E}[\|\alpha h(\omega_{t})-h^{\star}\|_{\mathbb{F}}^{2}]\leq\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|_{\mathbb{F}}^{2}]\exp(-m\lambda^{2}t).$
We use the above results to show that by choosing $\alpha$ sufficiently large, we can ensure that $\tau=\infty$ with high probability since $\lambda Id\preceq Dh(\omega_{t})D^{T}h(\omega_{t})$ holds. In the following result and throughout the rest of the paper, we denote the conditional probability and expectation given the random initialization $\omega_{0}$ as $\mathbb{P}_{\omega}$ and $\mathbb{E}_{\omega}$.
@@
$\mathbb{P}(\|\omega_{t}-\omega_{0}\|_{2}>r)\leq\frac{1}{\alpha r}\left(\frac{\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)}{m^{\frac{3}{2}}\lambda^{2}}+\frac{\mathrm{Lip}(\nabla\ell)}{m\lambda}{}\right)\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}_{\omega}[\|h^{\star}\|^{2}_{2}]}\lesssim\frac{1}{\alpha}.$ (9)
Proof of Theorem 2.
To prove the above statement, we will first use equation 1 and rewrite it to get an upper bound on the expected distance of $\omega_{t}$ from initialization. The resulting terms can then be further bounded by using the regularity of the loss function from Assumption 1 and the results from Theorem 1. We connect the expected distance to the probability of exiting a ball around the initialization using Markov’s inequality by $\mathbb{P}(\|\omega_{t}-\omega_{0}\|>\epsilon)\leq\frac{\mathbb{E}_{\omega}[\|\omega_{t}-\omega_{0}\|]}{\epsilon}$. Using equation 1, we can rewrite the distance of $\omega_{t}$ to the origin
$\leq\frac{1}{\alpha}\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{\mathbb{F}}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt+\frac{\sqrt{\eta_{\alpha}}}{\alpha}\|\int_{0}^{t}\sigma_{\alpha}(\omega_{t})dB_{t}\|_{2}$
$\mathbb{E}_{\omega}[\|\omega_{t}-\omega_{0}\|]$ $\leq\frac{\mathbb{E}_{\omega}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{\mathbb{F}}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{\alpha}+\frac{\sqrt{\eta_{\alpha}}\mathbb{E}_{\omega}[\|\int_{0}^{t}\sigma_{\alpha}(\omega_{t})dB_{t}\|_{2}]}{\alpha}$
$\leq\frac{\mathbb{E}_{\omega}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{\mathbb{F}}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{\alpha}+\frac{\left[\mathbb{E}_{\omega}[\int_{0}^{t}\mathrm{Tr}(\Sigma_{\alpha}(\omega_{s}))ds]\right]^{\frac{1}{2}}}{\alpha}.$
We will first apply Markov’s inequality to obtain
$\mathbb{P}(\|\omega_{t}-\omega_{0}\|_{2}\geq r)\leq\frac{\mathbb{E}_{\omega}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{\mathbb{F}}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{r\alpha}+\frac{\left[\mathbb{E}_{\omega}[\int_{0}^{t}\mathrm{Tr}(\Sigma_{\alpha}(\omega_{s}))ds]\right]^{\frac{1}{2}}}{r\alpha}.$ (10)
We will continue by first bounding the first part of the right-hand side in Equation 10.
$\frac{\mathbb{E}_{\omega}[\int_{0}^{t}\|D^{T}h(\omega_{t})\|_{\mathbb{F}}\|\nabla R(\alpha h(\omega_{t}))\|_{2}dt]}{r\alpha}$ $\overset{t\leq\tilde{T}}{\leq}\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)\mathbb{E}_{\omega}[\int_{0}^{t}\|\alpha h(\omega_{t})-h^{\star}\|_{2}dt]$
$=\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)\int_{0}^{t}\sqrt{\mathbb{E}_{\omega}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}]^{2}}dt$
$\overset{(\spadesuit)}{\leq}\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)\int_{0}^{t}\sqrt{\mathbb{E}_{\omega}[\|\alpha h(\omega_{t})-h^{\star}\|_{2}^{2}]}dt$
$\overset{(\clubsuit)}{\leq}\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)\int_{0}^{t}\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]\exp(-2m\lambda^{2}t)}dt$
$=\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]}\int_{0}^{t}\exp(-m\lambda^{2}t)dt$
$\leq\frac{1}{\alpha r}\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)\frac{\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|^{2}_{2}]}}{m^{\frac{3}{2}}\lambda^{2}},$
where $(\spadesuit)$ follows from Jensen’s inequality and $(\clubsuit)$ follows from Corollary 3. Next we will bound the second therm from the right-hand side of 10. We start by bounding
$\int_{0}^{t}Tr(\Sigma_{\alpha}(\omega_{s}))ds$ $=\int_{0}^{t}Tr\left(\mathbb{E}_{x}[\nabla\ell(x,\alpha h(\omega_{s})\nabla^{T}\ell(x,\alpha h(\omega_{s}))]-R(\alpha h(\omega_{s})R^{T}(\alpha h(\omega_{s})\right)ds$
$\leq\int_{0}^{t}Tr\left(\mathbb{E}_{x}[\nabla\ell(x,\alpha h(\omega_{s})\nabla^{T}\ell(x,\alpha h(\omega_{s}))]\right)ds$
$=\int_{0}^{t}\mathbb{E}_{x}[\|\nabla\ell(x,\alpha h(\omega_{s}))\|_{2}^{2}]ds$
$\overset{(\diamondsuit)}{\leq}\int_{0}^{t}\mathrm{Lip}(\nabla\ell)^{2}\mathbb{E}_{x}[\|\alpha h(\omega_{s})-h^{\star}\|_{\mathbb{F}}^{2}]ds$
$\overset{(\clubsuit)}{\leq}\frac{\mathrm{Lip}(\nabla\ell)^{2}\mathrm{Lip}(\nabla R)}{m}\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|_{\mathbb{F}}^{2}]\int_{0}^{t}\exp(-m\lambda^{2}s)ds,$
where we used the smoothness of $\ell$ in $\diamondsuit$ and Corollary 3 in $(\clubsuit)$. We can calculate that the value of $\int_{0}^{t}\exp(-m\lambda^{2}s)ds\leq\frac{1}{m\lambda^{2}}$, which gives us the following upper bound
$\frac{\sqrt{\mathbb{E}_{\omega}\left[\int_{0}^{t}Tr(\Sigma_{\alpha}(\omega_{s}))ds\right]}}{r\alpha}\leq\frac{\mathrm{Lip}(\nabla\ell)}{r\alpha m\lambda}\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}_{\omega}[\|\alpha h(\omega_{0})-h^{\star}\|_{\mathbb{F}}^{2}]}.$
By choosing an initialization such that $h(\omega_{0})=0$, we then get
$\mathbb{P}(\|\omega_{t}-\omega_{0}\|_{2}>r)\leq\frac{1}{\alpha r}\left(\frac{\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)}{m^{\frac{3}{2}}\lambda^{2}}+\frac{\mathrm{Lip}(\nabla\ell)}{m\lambda}{}\right)\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}_{\omega}[\|h^{\star}\|^{2}_{2}]}.$
In the following corollary, we will show that by choosing $\alpha$ appropriately, $\mathbb{P}(\tau<T)$ is arbitrarily small. This implies that we can effectively use our non-asymptotic bound that we established in Theorem 1 and Corollary 3 to get an upper bound on the training error with arbitrary large probability (depending on $\alpha$).
Corollary 4.
Under the same assumptions as Theorem 2, it follows for any $T>0$
$\mathbb{P}(\tau<T)\leq\mathcal{O}(\alpha^{-1}).$ (11)
Proof of Corollary 4.
In the following let $\tilde{\tau}:=\min\{\tau,T\}$. Since the right-hand side of equation 9 does not depend on $t$, we get
$\mathbb{P}(\tau<T)$ $=\mathbb{P}(\tilde{\tau}<T)$
@@
$\leq\mathcal{O}(\alpha^{-1}),$
where we used Theorem 2
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t}))]\leq\exp(-\lambda^{2}mt)+\mathcal{O}(\alpha^{-\frac{1}{q}}),$
for any $q>1$ and $t\leq T$ for any $T>0$.
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t}))]=\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau\geq T\}}]+\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<T\}}].$
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau\geq T\}}]\leq\exp(-\lambda^{2}mt),$
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<T\}}]\leq\mathcal{O}(\alpha^{-\frac{1}{q}}).$
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})\mathds{1}_{\{\tau<T\}}]$ $\leq\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t}))^{p}]^{\frac{1}{p}}\mathbb{E}_{\omega}[\mathds{1}_{\{\tau<T\}}^{q}]^{\frac{1}{q}}$ (12)
$\overset{\spadesuit}{=}\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})^{p}]^{\frac{1}{p}}\mathbb{E}_{\omega}[\mathds{1}_{\{\tau<T\}}]^{\frac{1}{q}}$ (13)
$=\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})^{p}]^{\frac{1}{p}}\mathbb{P}(\tau<T)^{\frac{1}{q}}$ (14)
$\overset{\clubsuit}{\leq}\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t})^{p}]^{\frac{1}{p}}\mathcal{O}(\alpha^{-\frac{1}{q}}),$ (15)
@@
$\mathbb{E}_{\omega}[\bar{R}(\alpha h(\omega_{t}))^{p}]\leq\bar{R}(\alpha h(\omega_{0}))^{p}e^{p}\mathbb{E}_{\omega}[\mathcal{E}_{t}]=\bar{R}(\alpha h(\omega_{0}))^{p}e^{p}.$
Plugging this result into equation 15 concludes the proof. ∎
@@
$\mathbb{E}_{\omega}[\|\alpha\bar{h}(\bar{\omega}_{t})-h^{\star}\|_{2}^{2}]\leq\frac{\mathrm{Lip}(\nabla R)}{m}\|h^{\star}\|^{2}_{2}\exp(-m\lambda^{2}t),$ (16)
@@
$\mathbb{P}(\|\bar{\omega}_{t}-\omega_{0}\|_{2}>r)\leq\frac{1}{\alpha r}\left(\frac{\|Dh(\omega_{0})\|_{\mathbb{F}}\mathrm{Lip}(\nabla R)}{m^{\frac{3}{2}}\lambda^{2}}+\frac{\mathrm{Lip}(\nabla\ell)}{m\lambda}{}\right)\sqrt{\mathrm{Lip}(\nabla R)\mathbb{E}_{\omega}[\|h^{\star}\|^{2}_{2}]}=\mathcal{O}\left(\alpha^{-1}\right).$
@@
$\mathbb{E}_{\omega}[\|\alpha\bar{h}(\bar{\omega}_{t})-\alpha h(\omega_{t})\|_{\mathbb{F}}]\leq 2\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}}\|h^{\star}\|_{\mathbb{F}}\exp(-m\lambda^{2}t).$
@@
$\mathbb{E}_{\omega}[\|\alpha\bar{h}(\bar{\omega}_{t})-\alpha h(\omega_{t})\|_{\mathbb{F}}]$ $\leq\mathbb{E}_{\omega}[\|\alpha\bar{h}(\bar{\omega}_{t})-h^{\star}\|_{\mathbb{F}}]+\mathbb{E}_{\omega}[\|\alpha h(\omega_{t})-h^{\star}\|_{\mathbb{F}}]$
$\overset{(\clubsuit)}{\leq}\sqrt{\mathbb{E}_{\omega}[\|\alpha\bar{h}(\bar{\omega}_{t})-h^{\star}\|_{\mathbb{F}}^{2}]}+\sqrt{\mathbb{E}_{\omega}[\|\alpha h(\omega_{t})-h^{\star}\|_{\mathbb{F}}^{2}]}$
$\overset{(\spadesuit)}{\leq}2\sqrt{\frac{\mathrm{Lip}(\nabla R)}{m}}\|h^{\star}\|_{\mathbb{F}}\exp(-m\lambda^{2}t),$